(6.4.7) Assume that females have pulse rates that are normally distributed with a mean of mu=76.0 beats per minute and a standard deviation of sigma=12.5 beats per minute.
a. If 1 adult female is randomly​ selected, find the probability that her pulse rate is between 72 beats per minute and 80 beats per minute.
The probability is
1
Expert's answer
2020-12-17T08:31:41-0500

Let "X=" pulse rate: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=76.0, \sigma=12.5"


"P(72<X<80)=P(X<80)-P(X\leq72)"

"=P(Z<\dfrac{80-76}{12.5})-P(Z\leq\dfrac{72-76}{12.5})"

"=P(Z<0.32)-P(Z\leq-0.32)\approx"

"\approx0.625516-0.374484=0.251032"

The probability is "0.251032."